OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –5) and is perpendicular to the line y = 3x – 6?

OpenStudy (anonymous):

Perpendicular lines have slopes that when multiplied equal to -1. So, what is the slope of the new line?

OpenStudy (anonymous):

x+3y+12 = 0

OpenStudy (anonymous):

In slope intercept form it is y= -x/3-4

OpenStudy (anonymous):

Product of slopes of mutually perpendicular line is -1. The given line has slope 3 and so 3m=-1 implies required slope is -1/3 Now equation line in slope-intercept form is y=mx+c Substitute m here y=-x/3+c. To get c substitute the point.....-5=-3/3+c implies c=-4 Then equation is y=-1/3x-4